# Open Problems, and Future Trends

### v1

@@ II-E2 Summary
AI technologies in the Industrial Metaverse generally require meeting requirements such as real-time capability[85], scalability[86], and interpretability[87]. However, there is still a significant research gap in interpretability. Many AI-assisted services and applications in the Metaverse rely on AI agents that are driven by black-box ML models, lacking explanation and interpretability capabilities. Developers, virtual world designers, and users of the Industrial Metaverse cannot understand the AI decision-making process from a mechanistic perspective. The lack of interpretability may lead to potential risks, as even a 1% error rate in decision-making can be unacceptable in industrial manufacturing. This impact not only affects economic benefits but also safety concerns.
@@ III Existing issues in Industrial Metaverse
In this section, we will delve into the crucial issues that currently exist in the Industrial Metaverse. With the rising complexity of connections and interactions between systems and devices, security threats and vulnerabilities have become a more pressing concern. Furthermore, the vast number of sensors, devices, and data involved in the Industrial Metaverse require efficient data management to prevent privacy breaches during data transmission[88]. Given that most enterprises prioritize profit maximization and face resource constraints, scientific resource optimization is necessary to reduce costs. We will elaborate on these challenges in the following sections. Fig. 8 illustrates the existing issues in Industrial Metaverse.
@@
III-A Security and Confidentiality
In the preceding segment, we presented an overview of how the Industrial Metaverse has attained extraordinary immersive encounters through the facilitation of cutting-edge technologies such as XR, DT, and AI, thereby establishing a novel paradigm for industrial administration. However, at the same time, it has also created unprecedented security threats. From data collection to the final decision-making stage, security and privacy challenges pervade the entire production cycle of the Industrial Metaverse. For example, the requirements for data integrity and availability, which are based on industrial data itself (i.e., ensuring that data flows from sensor collection endpoints to DT models without being maliciously modified or destroyed), as well as the requirements for authentication/authorization and traceability for employees (i.e., strict verification mechanisms for device operations and data access to ensure that only legitimate users can access them).
Although in the previous section, we mentioned the use of BC technology for anonymizing data and user identities through techniques such as identity verification and smart contracts, this alone cannot solve all security issues. In the following, we will organize existing solutions to security and privacy issues existing in various stages of the Industrial Metaverse based on the architecture diagram provided in Fig. 2 This includes data sources collected at the sensor end, as well as data and network security threats during subsequent cross-component transmission processes. We will also discuss security issues within components such as AI, BC, and XR, and further discuss possible research directions.
@@ III-A1 Sensor-side Industrial Data Source
When it comes to industrial source data, the emphasis lies in safeguarding the data itself. For instance, redundant information in visual content, which does not contribute to the training of models but is indiscriminately collected by sensors, can be obfuscated by overlaying a mask to mitigate potential privacy breaches. Validated operators can have their virtual images in the Metaverse replaced with generic mappings. Another approach involves employing black and white list mechanisms, where only specific information is retained, and any data outside the whitelist is immediately discarded and inaccessible. Wu et al.[89] propose a data misuse prevention mechanism called DAPter, wherein input data undergoes lightweight generator transformations to remove unnecessary information with minimal additional overhead. The blacklisting mechanism is commonly utilized for safeguarding user facial information. In this method, faces are decomposed into vector signals, comprising identity and attribute features. Subsequently, an anonymization model is generated to integrate and reconstruct the vectors based on the need for further processing[90]. This approach can also be employed for classifying features of industrial devices, such as device models and metric parameters. Additionally, traditional data encryption and storage schemes used in cloud-side end systems can be applied to Industrial Metaverse platforms. These include classical symmetric and asymmetric encryption techniques, cloud storage encryption, disk encryption, and more.
@@ III-A2 Data and Network Security Threats in Distributed Industrial Platforms
Another issue arising from the distributed architecture of the Industrial Metaverse is the potential for errors and attacks due to multiple untrusted but interdependent participating entities and the integration of non-intersecting data from multiple sources[102]. For instance, man-in-the-middle attacks or altering communication between users and servers may intercept/modify data packets, compromising the confidentiality, integrity, and availability of the system. Cross-site scripting attacks inject external malicious JavaScript code into websites, enabling attackers to inject malicious JavaScript code directly into client websites or have client websites request storage locations. Immersive interaction involves many devices such as wearable devices, headsets, base stations, and controllers, with large-scale data exchange. Despite manufacturers implementing encrypted communication, industrial data may still be accessed by adversaries through eavesdropping methods during transmission. Currently, efforts are underway to enhance industrial network security and anomaly detection. Kim et al.[103] utilize a system based on autoencoders and long short-term memory models to learn normal values of sensors and actuators for monitoring network attacks. Kwon et al.[104] study distributed network protocols widely used in distribution and transmission systems, proposing an anomaly detection method based on bidirectional recurrent networks and IEEE 1815.1. They verified the method with attack data from networked physical systems, successfully detecting eight types of attacks, including unauthorized command attacks and false data injection. However, existing detection systems deployed in Metaverse platforms often result in high computational costs and delays and are difficult to adapt to dynamic attack environments (such as atypical and polymorphic attacks). Therefore, network security researchers must further optimize and promote existing strategies to establish effective models.
@@ III-A3 Inherent Security Vulnerabilities of Enabling Technologies
Concerning AI technology, although Wang et al.[105] highlight its potential for detecting illicit entities and anomalous accounts within the Metaverse, such as discerning suspicious signals through correlating user-generated textual information and spatiotemporal activities, AI also poses various potential risks[106]. On one hand, the industrial platform interconnects a wide array of industrial devices, Metaverse users, collaborative enterprises, and applications, necessitating automated regulation and decision-making due to the intricacies of data management. This grants algorithms extensive operational privileges to reduce managerial costs. However, this approach may give rise to issues like biased decisions[107], low transparency[108], and operational fragility[109]. While certain existing work suggests hierarchical governance structures for algorithm accountability or assesses potential risks of algorithm engines based on AI explainability, a unified compliance governance protocol is yet to be established. On the other hand, due to the black-box nature of AI, malevolent attackers can infiltrate the enterprise source code with malicious code to cause system-level disruption. For instance, they may employ Generative Adversarial Networks (GANs) to fabricate false virtual data or embed malicious software during the model training process, which developers often find challenging to detect in early training results. Therefore, in the industrial context, modular interpretable AI may serve as a promising avenue for future research.
@@
BC records user information and production data in the Industrial Metaverse platform through a distributed ledger and facilitates anonymous P2P interactions through encryption, providing dependable data storage and transmission for enterprises and users[110]. Nevertheless, BC also harbors inherent security risks. Some emanate from the consensus mechanism, such as PoW, wherein an attacker with over half of the BC’s computing power can directly seize control of the entire platform system. Additionally, while consensus-based BC endows robust tamper resistance for industrial data, the computational overhead generated by existing mainstream consensus mechanisms is also a noteworthy concern. Another set of risks arises from BC’s key distribution mechanism. For instance, its decentralized key management implies that private keys might be concentrated and stored on specific entities or servers, rendering all keys vulnerable in the event of an attack on the entity. To avert private key leakage, multi-signature mechanisms can be employed, along with offline storage to mitigate network attacks.
Furthermore, XR-based Industrial Metaverse interaction technologies are susceptible to hacker attacks due to their numerous data interfaces. For example, various systems in factories utilize identity authentication methods such as fingerprint recognition and facial recognition for authorization. Hackers can gain access to device privileges by learning and reconstructing facial information. One immediate countermeasure is the design of novel authentication mechanisms. For instance, Wang et al.[111] devise an identity authentication scheme based on palmprints, which verifies through collecting the reverse scattered signal of acoustic signals, circumventing privacy leakage risks associated with image recognition. Duezguen et al.[112] leverage the concept of semantic connections within the ZeTA protocol to support identity authentication in XR context sharing. Finally, the authors evaluated the correlation between different interaction modes and risks. In the future, adopting compound multi-dimensional authentication technologies, coupled with timely firmware updates, may effectively mitigate privacy breaches during the interaction process. Lastly, beyond the aforementioned issues, addressing security challenges in the Industrial Metaverse necessitates keeping abreast of relevant legal documents and enterprise standards.
@@ III-B1 Computing Resource Allocation Issues
In collaborative multi-enterprise Industrial Metaverse platforms, some adaptable traditional distributed technologies are gradually being introduced. These include MEC and FL[114]. FL, particularly, is recognized for its privacy-preserving characteristics and excellent scalability as a distributed architecture that aggregates models remotely without exposing data. Al-Quraan et al.[115] discuss the potential applications of FL in various wireless networks, such as visible light communication and massive multiple-input multiple-output (mMIMO), to maintain network performance stability by training models that alleviate microbase station congestion. Nevertheless, despite offloading computation to the edge, FL may still face challenges in frequent information exchange between cloud servers and edge nodes due to the continuous generation of large volumes of new data in industrial production processes. Additionally, the increasing number of parameters and limited bandwidth resources may become bottlenecks for FL. Furthermore, due to differences in device models and performance in factories and varying requirements for model accuracy and computational resources in different scenarios, it is difficult to consider a universal FL aggregation design solution. Furthermore, the number of nodes in Industrial Metaverse systems is often indeterminable during application deployment, necessitating dynamic adjustments in resource allocation to accommodate the addition or removal of nodes[116]. In the face of extreme scenarios, the robustness of the deployment architecture must also be taken into consideration.
@@
Another common architecture is based on the allocation of computing resources in MEC. MEC, by deploying computing nodes near the terminals, is particularly suitable for high-concurrency industrial scenarios. Since edge computing nodes are close to physical devices, they can continue local data processing and decision-making even in cases of unstable or disconnected network connections[117]. Tan et al.[118] introduce DT into edge networks and designed power optimization based on Deep Q-Networks (DQN) for dynamic channel conditions in MEC. The proposed architecture was validated through simulations for applications in automated warehousing logistics and mobile production lines. In Industrial Metaverse platforms, VSPs are responsible for providing resources ordered by various manufacturers. However, excessive resource allocation may lead to resource over-provisioning issues. Ng et al.[119] propose a decision framework for VSPs in the Metaverse based on two-stage stochastic integer programming to minimize overall network costs considering the uncertainty of user demands. However, they did not consider solutions for resource demands in the case of large-scale resource requirements. To address these challenges, Chu et al.[120] decompose the Metaverse platform and utilized application similarity to improve resource utilization. They developed a self-learning framework based on Markov processes to capture implementation characteristics of the Metaverse. Considering the dynamic industrial environment, Yu et al.[121] propose an end-to-end self-organizing resource allocation framework under a 5G heterogeneous network, greatly enhancing Quality of Service (QoS).
@@
In scenarios involving multiple for-profit enterprises, the utility of VSPs in the Industrial Metaverse is considered a function of latency and resource payment costs induced by resource offloading. When the utility is non-negative, VSPs will be willing to invest more computational resources in enterprise collaboration[122]. One scheduling solution involves designing incentive mechanisms to maximize device utilization. Li et al.[123] introduce a distributed crowdsourcing framework called CrowdBC, where smart contracts are used for automated execution, and BC is used for user reputation management and incentive mechanisms. Reputation reflects the past performance of enterprise employees in task resolution. After submitting the execution plan, enterprise employees can request the system to perform task evaluation and record it. Luong et al.[122] propose using multi-agent auctions as an incentive mechanism to support semantic communication (SemCom) and computing resource allocation in the Metaverse. Transmitting semantic information instead of raw data can further reduce offloading latency. Additionally, as the key to synchronizing the real world with the virtual world lies in the data collected by IoT devices and sensors, Han et al.[124] design a dynamic resource allocation framework using a hybrid evolutionary dynamics approach to synchronize the Metaverse with IoT services and data. However, the aforementioned works did not consider the dynamic characteristics of each computing node in the Industrial Metaverse system, such as smart handheld terminals and AGVs. When nodes are predetermined, idle computational resources may remain underutilized when MEC handles computationally intensive tasks.

### v2

@@ II-F2 Summary
AI Technologies in the Industrial Metaverse typically need to meet requirements such as real-time performance[103], scalability[104], and interpretability[105]. Among these, the aspect of interpretability still presents a significant gap in research. In many AI-assisted services and applications within the Metaverse, decisions are made by AI agents, which are often driven by machine learning (ML) models functioning as black boxes, lacking the ability to explain their decisions. As a result, developers, virtual world designers, and users in the Industrial Metaverse may find it difficult to understand the mechanisms behind AI-driven decisions. This lack of interpretability poses potential risks; even a 1% decision-making error in industrial manufacturing can have unacceptable consequences, affecting not only economic outcomes but also the safety of production processes.
By developing more transparent AI models, operators’ trust in these systems can be enhanced, enabling more effective use of AI technology to support decision-making and operations. Such improvements contribute to creating a more efficient and safer collaborative environment within the Industrial Metaverse.
@@ III Existing issues in Industrial Metaverse
In this section, we will delve into the crucial issues that currently exist in the Industrial Metaverse. With the rising complexity of connections and interactions between systems and devices, security threats and vulnerabilities have become a more pressing concern. Furthermore, the vast number of sensors, devices, and data involved in the Industrial Metaverse require efficient data management to prevent privacy breaches during data transmission[106]. Given that most enterprises prioritize profit maximization and face resource constraints, scientific resource optimization is necessary to reduce costs. We will elaborate on these challenges in the following sections. Fig. 9 illustrates the existing issues in Industrial Metaverse. Table VI summarizes research efforts addressing existing challenges and issues in the Industrial Metaverse.
@@
III-A Security and Privacy
In the previous section, we outlined the immersive experiences and new management models brought by the Industrial Metaverse. However, this also introduces unprecedented security threats[107], with challenges related to security and privacy throughout the entire production cycle, from data collection to final decision-making. Specifically, the integrity and availability of industrial data (e.g., ensuring that data, from the sensor collection point to its input into the model, remains unaltered and undamaged) and the authentication, authorization, and traceability requirements for employees (i.e., ensuring that operations on devices and access to data go through a strict validation mechanism, allowing only authorized users to access them) are key issues that need to be addressed.
While some techniques mentioned in the previous section can anonymize data and user identities, they do not solve all security problems. We will organize existing solutions based on the architecture shown in Fig. 3. Then, we will discuss the security and privacy issues that may arise at different stages, including data collection, cross-component data transmission security threats, and internal security problems within components. Finally, we will explore potential research directions.
@@ III-A1 Sensor-side Industrial Data Source
When it comes to industrial source data, the emphasis lies in safeguarding the data itself. For instance, redundant information in visual content, which does not contribute to the training of models but is indiscriminately collected by sensors, can be obfuscated by overlaying a mask to mitigate potential privacy breaches. Validated operators can have their virtual images in the Metaverse replaced with generic mappings. Another approach involves employing black and white list mechanisms, where only specific information is retained, and any data outside the whitelist is immediately discarded and inaccessible. Wu et al.[108] propose a data misuse prevention mechanism called DAPter, wherein input data undergoes lightweight generator transformations to remove unnecessary information with minimal additional overhead. The blacklisting mechanism is commonly utilized for safeguarding user facial information. In this method, faces are decomposed into vector signals, comprising identity and attribute features. Subsequently, an anonymization model is generated to integrate and reconstruct the vectors based on the need for further processing. This approach can also be employed for classifying features of industrial devices, such as device models and metric parameters. Additionally, traditional data encryption and storage schemes used in cloud-side end systems can be applied to Industrial Metaverse platforms. These include classical symmetric and asymmetric encryption techniques, cloud storage encryption, disk encryption, and more.
@@ III-A2 Data and Network Security Threats in Distributed Industrial Platforms
Another issue arising from the distributed architecture of the Industrial Metaverse is the potential for errors and attacks due to multiple untrusted but interdependent participating entities and the integration of non-intersecting data from multiple sources. For instance, man-in-the-middle attacks or altering communication between users and servers may intercept/modify data packets, compromising the confidentiality, integrity, and availability of the system. Cross-site scripting attacks inject external malicious JavaScript code into websites, enabling attackers to inject malicious JavaScript code directly into client websites or have client websites request storage locations. Immersive interaction involves many devices such as wearable devices, headsets, base stations, and controllers, with large-scale data exchange. Despite manufacturers implementing encrypted communication, industrial data may still be accessed by adversaries through eavesdropping methods during transmission. Currently, efforts are underway to enhance industrial network security and anomaly detection. Kim et al.[120] utilize a system based on autoencoders and long short-term memory models to learn normal values of sensors and actuators for monitoring network attacks. Kwon et al.[121] study distributed network protocols widely used in distribution and transmission systems, proposing an anomaly detection method based on bidirectional recurrent networks and IEEE 1815.1. They verified the method with attack data from networked physical systems, successfully detecting eight types of attacks, including unauthorized command attacks and false data injection. However, existing detection systems deployed in Metaverse platforms often result in high computational costs and delays and are difficult to adapt to dynamic attack environments (such as atypical and polymorphic attacks). Therefore, network security researchers must further optimize and promote existing strategies to establish effective models.
@@ III-A3 Inherent Security Vulnerabilities of Enabling Technologies
Concerning AI technology, although Wang et al.[122] highlight its potential for detecting illicit entities and anomalous accounts within the Metaverse, such as discerning suspicious signals through correlating user-generated textual information and spatiotemporal activities, AI also poses various potential risks[123]. On one hand, the industrial platform interconnects a wide array of industrial devices, Metaverse users, collaborative enterprises, and applications, necessitating automated regulation and decision-making due to the intricacies of data management. This grants algorithms extensive operational privileges to reduce managerial costs. However, this approach may give rise to issues like biased decisions[124], low transparency[125], and operational fragility[126]. While certain existing work suggests hierarchical governance structures for algorithm accountability or assesses potential risks of algorithm engines based on AI explainability, a unified compliance governance protocol is yet to be established. On the other hand, due to the black-box nature of AI, malevolent attackers can infiltrate the enterprise source code with malicious code to cause system-level disruption. For instance, they may employ GANs to fabricate false virtual data or embed malicious software during the model training process, which developers often find challenging to detect in early training results. Therefore, in the industrial context, modular interpretable AI may serve as a promising avenue for future research.
@@
In the application of distributed ledger technology, the Industrial Metaverse platform records user information and production data, and enables anonymous P2P interactions through encryption methods, providing reliable data storage and transmission for enterprises and users. However, this also presents security risks, some of which stem from the consensus mechanism, such as PoW. If an attacker’s computational power exceeds half of the total power in the BC, they can take control of the entire platform system. Additionally, while consensus mechanisms can provide stronger tamper-proof capabilities for industrial data, the computational overhead generated by existing mainstream consensus mechanisms is also a significant concern[127]. Another risk arises from the key distribution mechanism, as decentralized key management means that private keys may be concentrated in an entity or server, and if that entity is attacked, all keys could be compromised. Currently, multi-signature mechanisms can be used to prevent private key leakage, and offline storage can help protect against network attacks.
In wireless communications, although 5G/6G wireless systems supported by the Metaverse bring many advantages, they also face security challenges. For example, Muhammad Adil et al.[128] point out that industrial automation supported by 5G/6G can operate on an independent platform that supports multiple interconnected applications performing different tasks. This deployment creates a metaspace at the network edge that faces potential threats such as DoS attacks, malware attacks, and authentication attacks. (Metaspace represents a logical space that manages physical space through control-plane separation, similar to software-defined networking and other enabling technologies[129]). Among these threats, malware can infiltrate the metaspace through unauthorized access and execute malicious activities. To mitigate these issues, it is best to use a robust yet lightweight authentication protocol that ensures the security of both customers and manufacturers. However, as previously mentioned, while various authentication schemes have been used to address the security of edge and cloud devices in the Industrial Metaverse, they are computationally complex and communication-intensive. Designing a lightweight authentication and data preservation scheme in the context of dynamic and hierarchical models may establish secure boundaries for subnets economically and efficiently. Additionally, Trusted Execution Environment (TEE) can serve as another promising security technology to address data confidentiality, availability, and integrity issues[130].
Using VR was once considered relatively secure because VR devices have a closed structure that prevents external attackers from directly observing displayed content. Unfortunately, recent researchers have emphasized that virtual reality users also face serious privacy threats when confronted with side-channel attacks[131]. On one hand, internal side-channel attackers can utilize motion sensor measurements to perform identity inference. For example, in XR systems, head movements play a crucial role in user interactions and navigation within virtual environments. Attackers can use machine learning techniques to analyze head movement data, associating specific movements with certain tasks or interactions, thereby reconstructing the algorithms being used and revealing sensitive information about product design and functionality. On the other hand, external attackers can exploit fluctuations in wireless signals emitted by VR devices to monitor password inputs[132]. For instance, various systems in factories rely on identity authentication, including fingerprint recognition, facial recognition, and more. Hackers can potentially learn and reconstruct facial information to gain access to devices. Currently, the most direct countermeasure is to design new authentication mechanisms. For example, Wang et al.[133] design a palmprint-based authentication scheme that collects backscattered acoustic signals for verification, thereby mitigating privacy risks associated with image recognition. In the future, adopting hybrid multidimensional authentication technologies, supplemented by timely firmware updates, may effectively counteract privacy breaches during interactions.
Lastly, beyond the aforementioned issues, addressing security challenges in the Industrial Metaverse necessitates keeping abreast of relevant legal documents and enterprise standards.
III-A4 Summary
In the above discussion, we have explored data and network security issues in the Industrial Metaverse. The security of industrial data sources at the sensor level requires mechanisms such as data masking and whitelist/blacklist strategies to reduce unnecessary data collection and mitigate privacy risks. To address malicious injection of low-quality data, solutions involving sensor node trustworthiness tracing and anomaly detection were proposed. Regarding data and network security threats in distributed industrial platforms, the decentralized architecture introduces risks of data breaches and network attacks[134]. BC technology and cloud-edge collaborative strategies can enhance data management and storage security, though challenges such as high energy consumption, information overload, and the “black box” effect remain. Finally, enabling technologies such as AI and XR also present potential security risks, including AI algorithm transparency issues and malicious exploitation, BC key management risks, and identity authentication issues in XR interactions. Addressing these challenges requires multi-dimensional technical solutions as well as supportive laws and standards.
@@ III-B1 Computing Resource Allocation Issues
In collaborative multi-enterprise Industrial Metaverse platforms, some adaptable traditional distributed technologies are gradually being introduced. These include MEC and FL[136]. Al-Quraan et al.[137] discuss the potential applications of FL in various wireless networks, such as visible light communication and massive multiple-input multiple-output (mMIMO), to maintain network performance stability by training models that alleviate microbase station congestion. Nevertheless, despite offloading computation to the edge, FL may still face challenges in frequent information exchange between cloud servers and edge nodes due to the continuous generation of large volumes of new data in industrial production processes. Additionally, the increasing number of parameters and limited bandwidth resources may become bottlenecks for FL. Furthermore, due to differences in device models and performance in factories and varying requirements for model accuracy and computational resources in different scenarios, it is difficult to consider a universal FL aggregation design solution[138]. Furthermore, the number of nodes in Industrial Metaverse systems is often indeterminable during application deployment, necessitating dynamic adjustments in resource allocation to accommodate the addition or removal of nodes[31]. In the face of extreme scenarios, the robustness of the deployment architecture must also be taken into consideration.
@@
Another common architecture is based on the allocation of computing resources in MEC[139]. MEC, by deploying computing nodes near the terminals, is particularly suitable for high-concurrency industrial scenarios. Since edge computing nodes are close to physical devices, they can continue local data processing and decision-making even in cases of unstable or disconnected network connections[140]. Tan et al.[141] introduce DT into edge networks and designed power optimization based on Deep Q-Networks (DQN) for dynamic channel conditions in MEC. The proposed architecture was validated through simulations for applications in automated warehousing logistics and mobile production lines. In Industrial Metaverse platforms, VSPs are responsible for providing resources ordered by various manufacturers. However, excessive resource allocation may lead to resource over-provisioning issues. Ng et al.[142] propose a decision framework for VSPs in the Metaverse based on two-stage stochastic integer programming to minimize overall network costs considering the uncertainty of user demands. However, they did not consider solutions for resource demands in the case of large-scale resource requirements. To address these challenges, Chu et al.[143] decompose the Metaverse platform and utilized application similarity to improve resource utilization. They developed a self-learning framework based on Markov processes to capture implementation characteristics of the Metaverse. Considering the dynamic industrial environment, Yu et al.[144] propose an end-to-end self-organizing resource allocation framework under a 5G heterogeneous network, greatly enhancing QoS.
@@
In scenarios involving multiple for-profit enterprises, the utility of VSPs in the Industrial Metaverse is considered a function of latency and resource payment costs induced by resource offloading. When the utility is non-negative, VSPs will be willing to invest more computational resources in enterprise collaboration. One scheduling solution involves designing incentive mechanisms to maximize device utilization. Li et al.[145] introduce a distributed crowdsourcing framework called CrowdBC, where smart contracts are used for automated execution, and BC is used for user reputation management and incentive mechanisms. Reputation reflects the past performance of enterprise employees in task resolution. After submitting the execution plan, enterprise employees can request the system to perform task evaluation and record it. As the key to synchronizing the real world with the virtual world lies in the data collected by IoT devices and sensors, Han et al.[146] design a dynamic resource allocation framework using a hybrid evolutionary dynamics approach to synchronize the Metaverse with IoT services and data. However, the aforementioned works did not consider the dynamic characteristics of each computing node in the Industrial Metaverse system, such as smart handheld terminals and AGVs. When nodes are predetermined, idle computational resources may remain underutilized when MEC handles computationally intensive tasks[147].